# Learning nuclear cross sections across the chart of nuclides with graph neural networks

### v2

@@ Abstract
In this work, we explore the use of deep learning techniques to learn how nuclear cross-sections change as we add or remove protons and neutrons. As a proof of principle, we focus on the neutron-induced reactions in the fast energy regime that are the most important in nuclear science and engineering. We use variational autoencoders (VAEs) and implicit neural representations (INRs) to build a learned feature representation space of nuclear cross sections and reduce the dimensionality of the problem. We then train graph neural networks (GNNs) on the resulting latent space to leverage the topological information encoded in the chart of isotopes and to capture the variations of cross sections with the number of particles in nuclei. We find that hypernetworks based on INRs significantly outperforms VAEs in encoding nuclear cross-sections. This superiority is attributed to INR’s ability to model complex, varying frequency details, which enables lower prediction errors when combined with GNNs. We also observe that GNN optimization is much more successful when performed in the latent space, whether using INRs or VAEs. However VAEs’ continuous representation also allows for direct GNN training in the original input space. We leverage these representational learning techniques and successfully predict cross sections for a $17\times 17$ block of nuclei with high accuracy and precision. Using clustering algorithms, we map groups of latent vectors into regions of the nuclear chart and show that VAEs and INRs can discover some of the neutron magic numbers. These findings suggest that deep-learning models based on representation encoding of cross-sections combined with graph neural networks hold significant potential in augmenting nuclear theory models, e.g., providing reliable estimates of covariances of cross sections, including cross-material covariances.
preprint: APS/123-QED
@@ I Introduction
Nuclear cross sections encode the probability of a subatomic particle to interact with an atomic nucleus, defined by its proton and neutron numbers $(Z,N)$ [1]. Each atomic nucleus can be involved in many different types of nuclear reactions often simply denoted as (incoming process, outgoing process), also called “channels”. For example, the $(n,\gamma)$ channel involves the absorption of a neutron and the emission of one or several photons ($\gamma$ rays) while the $(n,f)$ channel involves the absorption of a neutron resulting in the fission of the nucleus. Nuclear reaction libraries are formed by the set of all known cross sections for all known nuclei [2, 3, 4]. Such libraries are essential inputs to nuclear technology applications [5], which include energy production [6], nuclear medicine [7], and nuclear forensics [8]. They also play a critical role in our attempts at understanding the formation of elements in the universe [9, 10]. The importance of such libraries, and of nuclear data in general, in scientific research, was recently underlined in the 2023 Long Range Plan for Nuclear Science commissioned by the U.S. Department of Energy and the National Science Foundation [11].
@@
Measuring nuclear cross sections requires complex and expensive experimental facilities and is often impossible for short-lived radioactive nuclei that cannot be made as targets. While nuclear reaction theory has greatly improved over the past decades [12], its level of sophistication varies greatly depending on the nucleus and type of reaction. In very light nuclei, ab initio theory can successfully describe reactions at low energy in a fully quantum-mechanical framework [13]; in contrast, reactions of heavy nuclei with neutrons with just a few sub-eV energy already lead to the formation of a highly excited nucleus with an extremely complex structure that can only be described within a statistical framework [14, 15]. In practice, such statistical reaction theory relies on additional theoretical inputs describing nuclear properties such as the nuclear level density, the probability of capturing incident particles, the rate of decay through the emission of photons, neutrons, protons, etc. [16]. Today the best nuclear data libraries on the market largely rely on phenomenological models to determine such properties. These models are carefully calibrated to match realistic application requirements such as the criticality regimes of a nuclear reactor, at the price of a loss in generality and predictive power of the underlying theory [17, 4]. As a result, current models fail at consistently predicting the complex relationships between different types of cross sections in a single nucleus, or between cross sections in different nuclei across the chart of isotopes. This lack of predictive power is a significant impediment to efforts at designing new, cleaner or more efficient nuclear technologies. While research in fundamental reaction theory continues to progress, techniques originating from artificial intelligence and machine learning research offer an alternative viewpoint [18, 19, 20, 21].
@@
Indeed, each single cross section is a continuous one-dimensional function of the energy of the incident particle akin to a time sequence labeled by the number of protons $Z$ and neutrons $N$. In neutron-induced reactions in the fast regime – when the energy $E$ of the incident neutron is $E>100$ keV – cross sections are rather smooth functions of neutron energy (notwithstanding threshold effects). Therefore, dimensionality reduction techniques could in principle be applied to compress information and learn from the data the physically relevant latent space that underpin the cross sections.
@@
In addition, cross sections of nuclei with close $(Z,N)$ numbers often tend to have similar forms suggesting a correlation with the position on the 2D grid structure spanned by $Z$ and $N$. This observation suggests employing geometric deep learning for learning meaningful representations in the latent space, a technique that has recently emerged as an impactful sub-field adept at learning non-Euclidean sub-spaces [22]. Graph Neural Networks (GNNs) [23] in particular, provide an easy way to perform node-level, edge-level, and graph-level prediction tasks while being permutation invariant for structured grid-like data in such non-Euclidean sub-spaces, as has been shown for particle reconstruction [24].
@@
The goal of this paper is to expose the intricate relationships between cross sections of different nuclei by leveraging such state-of-the-art deep learning models. To this end, we employ two complementary techniques in succession. As the first step, we use either a variational autoencoder (VAE) [25] or an implicit neural representation (INR) to uncover the latent space of nuclear cross sections and reduce the dimensionality of the problem. In the second step, we use GNNs operating on such latent spaces to predict how cross sections transform when adding a neutron or a proton to a nucleus $(Z,N)$ in order to account for the implicit 2D geometric nature of the nuclear chart. The resulting network provides an efficient, consistent framework to predict unknown cross sections not too far from the training region.
@@
This paper is organized as follows: Section II gives a broad overview of the challenges facing nuclear theory when computing nuclear data. In Section III, we summarize the main aspects of our theoretical approach including the architecture of the network and the characteristics of the datasets used for training. Section IV shows the results for different GNN architectures and encoder approaches. Section V discusses our results in the context of related work on nuclear data before we summarize our findings in Section VI.
@@ II Theoretical models of nuclear data: Challenges
As mentioned in the introduction, nuclear reaction libraries are formed by the set of all known cross sections for all known nuclei. Most of this data is not simply a set of experimental data points but the result of an evaluation of all available data by subject matter experts. This evaluation is itself the outcome of a complex, multi-stage, iterative process that involves: (i) carefully analyzing available experimental results (vetting outliers, estimating uncertainties, etc.) (ii) adjusting model parameters of nuclear reaction codes such as TALYS [26] or EMPIRE [27] in such a way that the cross sections computed with such codes “match”, or are close to, experimental measurements (iii) processing the data (energy binning, heating of cross sections to room temperature values, etc.) (iv) verifying the nuclear reaction library through benchmarks, e.g., running a suite of codes simulating critical assemblies.
[…]
@@
In practice, the physics of a nuclear reaction is instead described by a suite of models with varying degrees of sophistication and phenomenology. These models play an essential role in the evaluation process and can be divided into several broad categories: (i) optical models describe the interaction of an incoming particle $a$ with a target $A$ (ii) level density models characterize the number of available energy levels $\rho(E,J,\pi)$ at a given excitation energy $E$, angular momentum $J$ and parity $\pi$ (iii) gamma-strength functions $\gamma_{SF}(E)$ characterize the probabilities of emitting photons as a function of energy $E$ (iv) fission transmission coefficients quantify the probability for a (heavy) nucleus to undergo fission as a function of energy, angular momentum and parity (v) pre-equilibrium reaction models describe interaction mechanisms between projectile and target that involve specific nuclear excited states.
@@
All these models are but different facets of the ideal quantum-mechanical evolution describe above. This implies that they should all be related to one another, at the very least indirectly. For example, let us assume one were to describe the structure of an atomic nucleus with a self-consistent mean-field model based on some effective nucleon-nucleon potential [28]. How this mean field is computed or parametrized should directly impact the optical potential (the real part of the optical potential does represent the nuclear mean field); the nuclear level density (through the excited states associated with the mean-field potential); the gamma-strength-function (which is typically related to functional derivatives of the mean field with respect to nuclear densities); fission transmission coefficients (which depend on how the mean-field potential changes with deformation), etc. Moreover, since the same effective nucleon-nucleon potential would be used to compute every single nucleus, nuclear properties would also be heavily correlated as a function of proton and neutron number.
@@
Therefore, even if such a theoretical model still contains parameters fitted to experimental data – in this case: the parameters of the effective nucleon-nucleon force, all nuclear properties are directly related to that single set of parameters. In a sensitivity analysis, one would find that variations of the model parameters result in correlated variations of the model outputs, i.e., the nuclear data properties of interest. In current approaches based on separate models for each type of nuclear data, such correlated variations, which manifest relationships between different data, are impossible to capture with nuclear theory alone.
@@
Instead, each nuclear model is treated as a separate black box independently parametrized by a handful of parameters. In a typical evaluation process, these parameters are adjusted in such a way that nuclear cross sections match experimental measurements and the resulting benchmarks are satisfied [3, 4, 29, 2]. These fits will often take the form of a functional dependency of model parameters in terms of proton and neutron number (or atomic mass number $A$), excitation energy or angular momentum. The nuclear reaction codes mentioned above typically implement a few different variants of such systematics. By construction, these systematics become highly questionable in nuclei or at energies that are beyond the range of their fit, which directly impacts the reliability of nuclear data libraries in these regions.
@@
By employing a collection of independent nuclear models locally calibrated to experimental data, it thus becomes nearly impossible to (1) make reliable predictions outside the range of currently available measurements (2) capture the relationships (or correlations) between nuclear data within the same nucleus, or between different nuclei. In this paper, we will show that deep-learning techniques provide an alternative solution to point (2) .
@@
Figure 1: Schematic representation of our framework. The initial data is a (set of) nuclear cross sections represented as a vector indexed by the energy of the incident particle. The data is encoded into a latent representation using either a variational autoencoder (VAE) or an implicit neural representation (INR) network (representation learning). A graph neural network (GNN) then learns how the data in the latent space in nucleus $(Z,N)$ transforms into neighboring nuclei in order to make predictions in nucleus $(Z^{\prime},N^{\prime})$. The decoder then transforms the predicted vector in the latent space back into the space of nuclear cross sections (latent space predictions).
@@ III Methodology
The limited number of available data for individual cross sections together with the potentially complex relationships between cross sections of different nuclei can make the training of deep learning models challenging. To overcome such a challenge, we separate the overall learning problem into two interconnected components, namely a representation learning task and a prediction task, as shown in Fig. 1. In the representation learning task, we are learning the low-dimensional, latent space representation of cross sections; in the prediction task, we make predictions to unknown nuclei from such latent representations and use the decoder trained in the first task to bring them back into the original space of cross sections. In contrast to the space of cross sections, the latent space provides a more compact yet meaningful representation that allows the model to learn more rapidly and generalize better. Additionally, the separation of these two steps ensures a clear delineation between the process of learning the representation, which may require data augmentation, i.e., interpolation and scaling of the cross-section data, and the actual prediction tasks, which should only involve the cross section data that physically exist, i.e., a cross section that is not augmented with additional, synthetic data or modified.
As schematically illustrated in Fig. 1(a), the representation learning component will focus on learning the underlying latent representation space of all known cross sections. In the following discussion, we note $\sigma_{\alpha}(E)$ the cross section of the nuclear reaction, or channel, $\alpha$ as a function of the energy $E$ of the incident particle, where the energy is discretized on a finite grid $E\rightarrow\{E_{o}\}_{o=1,m}$, where $m$ is the dimension of the original data vector. For the encoding of the cross-section, we train either a 1D-convolutional VAE [30] or an encoding-decoding hypernetwork [31] that predicts the weights for the INR model [32] on data from $s$ nuclei to learn a $k$-dimensional latent space for various cross sections. By design, $k<m$ so that the size of the feature matrix is reduced to $\mathbf{X}\in\mathbb{R}^{s\times k}$ while keeping the implicit grid intact.
@@ III.1 Dataset
Our dataset is a subset of the TENDL dataset of nuclear cross sections [29]. In practice, we only consider the $(n,el)$ (i.e., elastic), $(n,n^{\prime})$ (i.e., inelastic), $(n,\gamma)$ and $(n,2n)$ cross sections for all nuclei included in TENDL. Since cross sections are probabilities per unit area, their values are always positive: $\sigma_{\alpha}(E)\geq 0$ for all channels $\alpha$. The TENDL dataset contains $s=2151$ nuclei and can be represented in a 2D grid corresponding to the chart of isotopes and labeled as $(N,Z)$ with $N$ the neutron number and $Z$ the proton number of the nucleus. For the elastic, inelastic, $(n,\gamma)$ and $(n,2n)$ channels the original cross section is defined on an energy grid of $m=256$ points regularly spaced between 0 and 20 MeV. The values of each type of cross section are normalized by dividing by their maximum across the dataset. As discussed before, the data is then embedded into a lower $k$-dimensional latent space with $k=32$ using either a pre-trained VAE or INR model, resulting in a final data matrix of size $(s\times k)=(2151\times 32)$.
The size of the TENDL dataset is rather small to train deep learning models such as VAEs or INRs. This can be mitigated using data augmentation techniques. In the representation learning task, we only focus on the cross section data, i.e., describing curves, rather than any underlying physical relationship and properties. The goal is simply to construct a continuous, multi-dimensional latent space of curve shapes. Since we do not directly assign any physical properties to such a space, we can adopt an interpolation-based data augmentation strategy, in which we produce new training examples by linearly interpolating between two randomly picked cross section samples with weights that add up to 1.0. Essentially the latent space thus generated provides a super-set of all valid cross-sections while at the same time allowing for a smooth transition between them for better optimization in the cross-section prediction task with the GNN. For the latter, only actual cross sections in existing nuclei are used as training data.

### v3

@@ Abstract
In this work, we explore the use of deep learning techniques to learn how nuclear cross sections change as we add or remove protons and neutrons. As a proof of principle, we focus on the neutron-induced reactions in the fast energy regime. Our approach follows a two-stage learning framework. First, we apply representation learning to encode cross section data into a latent space using either variational autoencoders (VAEs) or implicit neural representations (INRs). Then, we train graph neural networks (GNNs) on the resulting embeddings to predict missing values across the nuclear chart by leveraging the topological structure of neighboring isotopes. We demonstrate accurate cross section predictions within a $9\times 9$ block of missing nuclei. We also find that the optimal GNN training strategy depends on the type of latent representation used, with VAE embeddings performing best under end-to-end optimization in the original space, while INR embeddings achieve better results when the GNN is trained only in the latent space. Furthermore, using clustering algorithms, we map groups of latent vectors into regions of the nuclear chart and show that VAEs and INRs can discover some of the neutron magic numbers. These findings suggest that deep-learning models based on the representation encoding of cross sections combined with graph neural networks holds significant potential in augmenting nuclear theory models, e.g., by providing reliable estimates of covariances of cross sections, including cross-material covariances.
@@ I Introduction
Nuclear cross sections encode the probability of a subatomic particle to interact with an atomic nucleus, defined by its proton and neutron numbers $(Z,N)$ [1]. Each atomic nucleus can be involved in many different types of nuclear reactions often simply denoted as (incoming process, outgoing process), also called “channels”. For example, the $(n,\gamma)$ channel involves the absorption of a neutron and the emission of one or several photons ($\gamma$ rays) while the $(n,f)$ channel involves the absorption of a neutron resulting in the fission of the nucleus. Nuclear reaction libraries are formed by the set of all known cross sections for all known nuclei [2, 3, 4]. Such libraries are essential inputs to nuclear technology applications [5], which include energy production [6], nuclear medicine [7], and nuclear forensics [8]. They also play a critical role in our attempts at understanding the formation of elements in the universe [9, 10]. The importance of such libraries, and of nuclear data in general, in scientific research was recently underlined in the 2023 Long Range Plan for Nuclear Science commissioned by the U.S. Department of Energy and the National Science Foundation [11].
@@
Measuring nuclear cross sections requires complex and expensive experimental facilities and is often impossible for short-lived radioactive nuclei that cannot be made into targets or created in radioactive beams. While nuclear reaction theory has greatly improved over the past decades [12], its level of sophistication varies greatly depending on the nucleus and type of reaction. In very light nuclei, ab initio theory can successfully describe reactions at low energy in a fully quantum-mechanical framework [13]; in contrast, reactions of heavy nuclei with neutrons with just a few sub-eV energy already lead to the formation of a highly excited nucleus with an extremely complex structure that can only be described within a statistical phenomenology framework [14, 15]. In practice, such statistical reaction theory relies on additional theoretical inputs describing nuclear properties such as the nuclear level density; the probability of capturing incident particles; the rate of decay through the emission of photons, neutrons, protons; etc. [16]. Today the best nuclear data libraries on the market largely rely on phenomenological models to determine such properties. These models are carefully calibrated to match realistic application requirements such as the criticality regimes of a nuclear reactor, at the price of a loss in generality and predictive power of the underlying theory [17, 4]. As a result, current models fail at consistently predicting the complex relationships between different types of cross sections in a single nucleus, or between cross sections in different nuclei across the chart of isotopes. This lack of predictive power is a significant impediment to efforts at designing new, cleaner, or more efficient nuclear technologies. While research in fundamental reaction theory continues to progress, techniques originating from artificial intelligence and machine learning research offer an alternative viewpoint [18, 19, 20, 21].
@@
Indeed, each single cross section is a continuous one-dimensional function of the energy of the incident particle (akin to a time sequence) labeled by the number of protons $Z$ and neutrons $N$. In neutron-induced reactions in the fast regime—when the energy $E$ of the incident neutron is $E>100$ keV— cross sections are rather smooth functions of neutron energy (notwithstanding threshold effects). Therefore, dimensionality reduction techniques could in principle be applied to compress information and learn from the data the physically relevant latent space that underpin the cross sections.
@@
In addition, cross sections of nuclei with close $(Z,N)$ numbers often tend to have similar forms, suggesting a correlation with the position on the 2D grid structure spanned by $Z$ and $N$. This observation suggests employing geometric deep learning for learning meaningful representations in the latent space, a technique that has recently emerged as an impactful sub-field adept at learning non-Euclidean sub-spaces [22]. Graph Neural Networks (GNNs) [23] in particular provide an easy way to perform node-level, edge-level, and graph-level prediction tasks while being permutation-invariant for structured grid-like data in such non-Euclidean sub-spaces, as has been shown for particle reconstruction [24].
@@
The goal of this paper is to effectively learn the intricate relationships between cross sections of different nuclei by leveraging such state-of-the-art deep learning models. To this end, we employ two complementary techniques in succession. As the first step, we use either a variational autoencoder (VAE) [25] or an implicit neural representation (INR) to uncover the latent space of nuclear cross sections and reduce the dimensionality of the problem. In the second step, we use GNNs to account for the implicit 2D geometric nature of the nuclear chart, where the graph neighborhood is defined by similar $Z$ and $N$ numbers, and the learned latent space are used as graph node features. The resulting network provides an efficient and consistent framework to predict unknown cross sections not too far from the training region.
@@
This paper is organized as follows: Section II gives a broad overview of the challenges facing nuclear theory when computing nuclear data. In Section III, we summarize the main aspects of our theoretical approach, including the architecture of the network and the characteristics of the datasets used for training. Section IV shows the results for different GNN architectures and encoder approaches. Section V discusses our results in the context of related work on nuclear data before we summarize our findings in Section VI.
@@ II Theoretical models of nuclear data: Challenges
As mentioned in the introduction, nuclear reaction libraries are formed by the set of all known cross sections for all known nuclei. Most of this data is not simply a set of experimental data points but the result of an evaluation of all available data by subject matter experts. This evaluation is itself the outcome of a complex, multi-stage, iterative process that involves: (i) carefully analyzing available experimental results (vetting outliers, estimating uncertainties, etc.); (ii) adjusting model parameters of nuclear reaction codes such as TALYS [26] or EMPIRE [27] in such a way that the cross sections computed with such codes “match”, or are close to, experimental measurements; (iii) processing the data (energy binning, heating of cross sections to room temperature values, etc.); and (iv) verifying the nuclear reaction library through benchmarks, e.g., running a suite of codes simulating critical assemblies.
From a physics perspective, nuclear reactions are extremely complex processes. Ideally, one would like to describe any reaction as a quantum-mechanical evolution from an initial state, e.g., an incoming particle $a$ with energy $E_{a}$, spin $J_{a}$, and parity $\pi_{a}$ impinging on a target at rest $A$ with energy $E_{A}$, spin $J_{A}$, and parity $\pi_{A}$ to a final state such as two reaction products $b$ and $B$ with respective energies $E_{b}$ ($E_{B}$), spins $J_{b}$ ($J_{B}$), and parities $\pi_{b}$ ($\pi_{B}$). Such a description would require the perfect characterization of the quantum-mechanical state of all reaction products $a$, $A$, $b$, and $B$ as well as the exact knowledge of the nuclear many-body Hamiltonian $\hat{H}$, since the unitary quantum-mechanical time evolution operator is $e^{-i\hat{H}t/\hbar}$. Both are out of reach of current nuclear theory.
@@
In practice, the physics of a nuclear reaction is instead described by a suite of models with varying degrees of sophistication and phenomenology. These models play an essential role in the evaluation process and can be divided into several broad categories: (i) optical models describe the interaction of an incoming particle $a$ with a target $A$; (ii) level density models characterize the number of available energy levels $\rho(E,J,\pi)$ at a given excitation energy $E$, angular momentum $J$, and parity $\pi$; (iii) gamma-strength functions $\gamma_{SF}(E)$ characterize the probabilities of emitting photons as a function of energy $E$; (iv) fission transmission coefficients quantify the probability for a (heavy) nucleus to undergo fission as a function of energy, angular momentum, and parity; and (v) pre-equilibrium reaction models describe interaction mechanisms between projectile and target that involve specific nuclear excited states.
@@
All these models are but different facets of the ideal quantum-mechanical evolution described above. This implies that they should all be related to one another, at the very least indirectly. For example, let us assume one were to describe the structure of an atomic nucleus with a self-consistent mean-field model based on some effective nucleon-nucleon potential [28]. How this mean field is computed or parametrized should directly impact the optical potential (the real part of the optical potential does represent the nuclear mean field), the nuclear level density (through the excited states associated with the mean-field potential), the gamma-strength-function (which is typically related to functional derivatives of the mean field with respect to nuclear densities), fission transmission coefficients (which depend on how the mean-field potential changes with deformation), etc. Moreover, since the same effective nucleon-nucleon potential would be used to compute every single nucleus, nuclear properties would also be heavily correlated as a function of proton and neutron number.
@@
Therefore, even if such a theoretical model still contains parameters fitted to experimental data—in this case, the parameters of the effective nucleon-nucleon force—all nuclear properties are directly related to that single set of parameters. In a sensitivity analysis, one would find that variations of the model parameters result in correlated variations of the model outputs, i.e., the nuclear data properties of interest. In current approaches based on separate models for each type of nuclear data, such correlated variations, which manifest relationships between different data, are impossible to capture with nuclear theory alone.
@@
Instead, each nuclear model is treated as a separate black box independently parametrized by a handful of parameters. In a typical evaluation process, these parameters are adjusted in such a way that nuclear cross sections match experimental measurements and the resulting benchmarks are satisfied [3, 4, 29, 2]. These fits will often take the form of a functional dependency of model parameters in terms of proton and neutron number (or atomic mass number $A$), excitation energy, or angular momentum. The nuclear reaction codes mentioned above typically implement a few different variants of such systematics. By construction, these systematics become highly questionable in nuclei or at energies that are beyond the range of their fit, which directly impacts the reliability of nuclear data libraries for extrapolations.
@@
By employing a collection of independent nuclear models locally calibrated to experimental data, it thus becomes nearly impossible to (1) make reliable predictions outside the range of currently available measurements and (2) capture the relationships (or correlations) between nuclear data within the same nucleus, or between different nuclei. In this paper, we will show that deep-learning techniques provide an alternative solution to both points.
@@
Figure 1: Schematic representation of our framework. The initial data is a (set of) nuclear cross sections represented as a vector indexed by the energy of the incident particle. Panel (a): The data is encoded into a latent representation using either a variational autoencoder (VAE) or an implicit neural representation (INR) network (representation learning). Panel (b): A graph neural network (GNN) then learns how the data in the latent space in nucleus $(Z,N)$ transforms into neighboring nuclei in order to make predictions in nucleus $(Z^{\prime},N^{\prime})$. The decoder then transforms the predicted vector in the latent space back into the space of nuclear cross sections (latent space predictions).
@@ III Methodology
The limited number of available data for individual cross sections together with the potentially complex relationships between cross sections of different nuclei can make the training of deep learning models challenging. To overcome such a challenge, we separate the overall learning problem into two interconnected components, namely a representation learning task and a prediction task, as shown in Fig. 1. In the representation learning task, we are learning the low-dimensional, latent space representation of cross sections; in the prediction task, we make predictions to unknown nuclei from such latent representations and use the decoder trained in the first task to bring them back into the original space of cross sections. In contrast to the space of cross sections, the latent space provides a more compact yet meaningful representation that allows the model to learn more rapidly and generalize better. Additionally, the separation of these two steps ensures a clear delineation between the process of learning the representation, which may require data augmentation, i.e., interpolation and scaling of the cross section data, and the actual prediction tasks, which should only involve the cross section data that physically exist, i.e., a cross section that is not augmented with additional, synthetic data or modified.
As schematically illustrated in Fig. 1(a), the representation learning component will focus on learning the underlying latent representation space of all known cross sections. In the following discussion, we note $\sigma_{\alpha}(E)$ the cross section of the nuclear reaction, or channel, $\alpha$ as a function of the energy $E$ of the incident particle, where the energy is discretized on a finite grid $E\rightarrow\{E_{o}\}_{o=1,m}$, where $m$ is the dimension of the original data vector. For the encoding of the cross section, we train either a 1D-convolutional VAE [30] or an encoding-decoding hypernetwork [31] that predicts the weights for the INR model [32] on data from $s$ nuclei to learn a $k$-dimensional latent space for various cross sections. By design, $k<m$ so that the size of the feature matrix is reduced to $\mathbf{X}\in\mathbb{R}^{s\times k}$ while keeping the implicit grid intact.
@@ III.1 Dataset
Our dataset is a subset of the TENDL dataset of nuclear cross sections [29]. In practice, we only consider the $(n,el)$ (i.e., elastic), $(n,n^{\prime})$ (i.e., inelastic), $(n,\gamma)$, and $(n,2n)$ cross sections for all nuclei included in TENDL. Since cross sections are probabilities per unit area, their values are always positive: $\sigma_{\alpha}(E)\geq 0$ for all channels $\alpha$. The TENDL dataset used contains $s=2171$ nuclei and can be represented in a 2D grid corresponding to the chart of isotopes and labeled as $(Z,N)$ with $Z$ the proton number and $N$ the neutron number of the nucleus. For the elastic, inelastic, $(n,\gamma)$, and $(n,2n)$ channels the original cross section is defined on an energy grid of $m=256$ points regularly spaced between 0 and 20 MeV. The values of each type of cross section are normalized by dividing by their maximum across the dataset prior to network training. After training, the data is then embedded into a lower $k$-dimensional latent space with $k=32$ using either a pre-trained VAE or INR model, resulting in a final data matrix of size $(s\times k)=(2171\times 32)$, which serve as input for the downstream prediction task.
The size of the TENDL dataset is rather small to train deep neural networks. This can be mitigated using data augmentation techniques. In the representation learning task, we only focus on the cross section data, i.e., describing curves, rather than any underlying physical relationship and properties. The goal is simply to construct a continuous, multi-dimensional latent space of curve shapes. Since we do not directly assign any physical properties to such a space, we can adopt an interpolation-based data augmentation strategy, in which we produce new training examples by linearly interpolating between two randomly picked cross section samples with weights that add up to 1.0. Essentially the latent space thus generated provides a super-set of all valid cross sections while at the same time allowing for a smooth transition between them for better optimization in the cross section prediction task with the GNN. For the latter, only actual cross sections in existing nuclei are used as training data. We found that this type of data augmentation improves the encoding performance of the VAE model, whereas the INR model does not benefit from it during training.